A. Compute the missing information to prepare the Balance sheet given the following data.

a. Assets:

  • Cash - 7,000.00
  • Accounts Receivable - 600.00
  • Inventory - 700.00
  • Equipment - 1,000.00
  • Total assets - ?

b. Liabilities:

  • Accounts Payable - 1,900.00
  • Loan Payable - ?
  • Total Liabilities - 5,400.00

c. Equity:

  • Owner's Equity - 2,700.00
  • Retained Earnings - 1,200.00
  • Total Equity - ?
1
Expert's answer
2021-11-14T17:32:27-0500

a)Total assets=7,000+600+700+1,000=9,300


b)Loans payable=5,400-1,900=3,500


c)Total equity=2,700+1,200=3,900